OpenStudy (anonymous):

Show cos 20 * cos 40 * cos 80 =1/8

OpenStudy (raden):

supposed the value of (cos 20 * cos 40 * cos 80 = x) multiply by sin20 to both sides, the equation above becomes sin20 * cos 20 * cos 40 * cos 80 = x sin20 look at that sin20cos20 = 1/2 sin(2(20)) = 1/2 sin40, now we have (sin20 * cos 20) * cos 40 * cos 80 = x sin20 1/2 sin40 * cos 40 * cos 80 = x sin20 and again by using double angle formula for sin, we can change sin40 * cos40 = 1/2 sin(2(40)) = 1/2 sin80, now we have 1/2 (sin40 * cos 40) * cos 80 = x sin20 1/2 * (1/2 sin80) * cos80 = xsin20 1/4 sin80 * cos80 = x sin20 and again again we can rwrite sin80 * cos80 becomes 1/2 sin(2(80)) = 1/2sin160 now we have : 1/4 (sin80 * cos80) = x sin20 1/4 * 1/2sin160 = x sin20 remember the identity : sin(180-x) = sinx therefore, sin160 = sin(180-20) = sin20 finally we have an equation : 1/4 * (1/2sin160) = x sin20 1/4 * 1/2 * sin20 = x sin20 1/8 sin20 = x sin20 divide by sin20 to both sides, finally we get 1/8 sin20/sin20 = x sin20/sin20 1/8 = x or 1/8 = cos 20 * cos 40 * cos 80 QED ^_^